"Peanut Plus" containing 50% peanuts and "Deluxe Nuts" containing 25% peanuts are to be combined to form 100 lbs. of mixed nuts that will contain 35% peanuts. How many pounds of each should be used? Write a system of equations for the problem and then solve to answer please. Thank you so much!
Answer by richwmiller(17219)

x+y=100
.5x+.25y=.35*100
x=40
y=60
check
.5*40+.25*60=.35*100
20+15=35
ok